OpenStudy (anonymous):

Jack has decided that he needs to get more exercise so he is building a running regimen for himself. He decided that, beginning Sunday, he'll run 10 feet. On each subsequent day, he'll run double the distance he ran the previous day. So he'll run 20 feet on Monday, 40 feet on Tuesday, and so on. How far will Jack have to run on the 14th day? If Jack sticks to his routine, how far will he have run in 14 days?

OpenStudy (dan815):

i mean geometric series

OpenStudy (anonymous):

thats what i thought

OpenStudy (anonymous):

i just dont know how to set up the formula

OpenStudy (dan815):

how many times will it have doubled by the 14th day

OpenStudy (dan815):

if its doubling everyday

OpenStudy (anonymous):

14

OpenStudy (anonymous):

or 13

OpenStudy (dan815):

ya 13

OpenStudy (anonymous):

so where does that go into the formula

OpenStudy (dan815):

what does it mean to double 13 times?

OpenStudy (dan815):

10*2*2*2*2....*2

OpenStudy (dan815):

10*2^13

OpenStudy (anonymous):

so 20^13

OpenStudy (dan815):

no

OpenStudy (dan815):

exponent before multiplication

OpenStudy (dan815):

its just 10 * 2*2*2*...*2, 10 multiplying 13 of 2s

OpenStudy (anonymous):

what about a formula?

OpenStudy (anonymous):

or is this easier

OpenStudy (dan815):

why dont u come up with that

OpenStudy (dan815):

look at the pattern

OpenStudy (anonymous):

81920

OpenStudy (dan815):

1st day = 10 2nd day = 10*2 3rd day= 10*2^2 . . >

OpenStudy (dan815):

nth day =?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

the common ratio is 20?

OpenStudy (dan815):

what?

OpenStudy (anonymous):

is the common ratio 20

OpenStudy (dan815):

what is that

OpenStudy (anonymous):

i dont know what it is good enough to explain to you haha

OpenStudy (anonymous):

i think it is 20 though hold on

OpenStudy (dan815):

no

OpenStudy (anonymous):

it is 2?

OpenStudy (dan815):

1st day = 10 2nd day = 10*2 3rd day= 10*2^2 4th day =10*2^3 .. . . nth day=10*2^(n-1)

OpenStudy (anonymous):

when do you use n-1 and n+1

OpenStudy (dan815):

look at the pattern

OpenStudy (dan815):

1st day = 10*2^0 2nd day = 10*2^1 3rd day = 10*2^2

OpenStudy (dan815):

10*2^(day number -1)

OpenStudy (anonymous):

but when do you know when to use n-1 or n+1

OpenStudy (anonymous):

why isnt it (n+1)

OpenStudy (dan815):

i feel like u are just tryna look for formulas

OpenStudy (dan815):

instead of just understanding it

OpenStudy (anonymous):

yeah i am :P

OpenStudy (anonymous):

i dont understand algebra

OpenStudy (dan815):

theres no formulas, u are gonna forget the formula

OpenStudy (dan815):

in a couple days

OpenStudy (dan815):

but u understand this u wont ever forget it

OpenStudy (anonymous):

when you do 10*2^(day number -1) why isnt it 10*2^(day number +1)

OpenStudy (dan815):

u tell me -_-

OpenStudy (dan815):

does it work with the pattern?

OpenStudy (dan815):

1st day = 10*2^(1+1)? or 10*2^(1-1)

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

so he will have to walk 81920 steps on the 14 day

OpenStudy (anonymous):

how do i find out how many steps he walked without adding each individual day?

OpenStudy (anonymous):

@Luigi0210